Let X be a random variable satisfying P(-1 X 1) = 0.3, P(X = 1.5) = 0.1, P(1.5 X P(3 X 7.4) 0.3, P(X 10)0.2 2) = 0.2 Find (i) P(X 2 1.3) (ii) P(X 2.3) ii P(1.5< X 2) (iv) P(1.5 3X 38)
Answer #1

informations given,

  • P(-1\leq X\leq 1)=0.3
  • PIX-1.5) 0.1
  • P(1.5\leq X \leq 2)=0.2
  • P(3\leq X \leq 7.4)=0.3
  • P(X=10)=0.2

a)

P(X\geq 1.3)

=P(1.5\leq X \leq 2)+P(3 \leq X\leq 7.4)+P(X=10)

[the interval  (1.5 < x 〈2) include the point X=1.5 , so we don't have to put P(X=1.5) in the above expression]

=0.2+0.3+0.2

=0.7

b)

P(X \leq 2.3)

=P(-1\leq X \leq 1)+P(1.5\leq X \leq 2)

[the interval  (1.5 < x 〈2) include the point X=1.5 , so we don't have to put P(X=1.5) in the above expression]

=0.3+0.2

=0.5

c)

P(1.5<X\leq 2)

=P(1.5\leq X\leq 2)-P(X=1.5)

=0.2-0.1

=0.1

d)

P(1.5\leq X \leq 8)

=P(1.5\leq X \leq 2)+P(13\leq X \leq 7.4)

=0.2+0.3

=0.5